Thomas buys a cardboard sheet that is 12 by 8 inches. Create an equation for this situation, find the zeros, and sketch the function.

Found 3 solutions by josgarithmetic, Riley12, ikleyn:
Answer by josgarithmetic(39614)

What more is given? All you have so far is cardboard sheet, length 12 inches, width 8 inches. What is to be done with this sheet?

Answer by Riley12(3)

the equation:
V= (96 - 40x + 4x^2) x = 96x - 40x^2 + 4x^3
V=(12-2x)(8-2x)(x)
The zeroes of x (12 - 2x) (8 -2x) = 0
are x=0, x=4 and x=6

Answer by ikleyn(52756)

.
Thomas buys a cardboard sheet that is 12 by 8 inches. Create an equation for this situation, find the zeros, and sketch the function.
~~~~~~~~~~~~


For WHICH situation ?



The post, formulated in this way,  DOES  NOT  deserve any consideration at this forum .

It  ONLY  deserves to be deleted or thrown to a  GARBAGE  BOX . . .


See the rules posting to this forum at this web-page

https://www.algebra.com/tutors/students/ask.mpl?action=ask_question&topic=Equations&return_url=http://www.algebra.com/algebra/homework/equations/

from which you post your problems.



It is assumed that you read these rules before posting.

It is also assumed that you do understand what is written in that page and follow the rules.